INSTRUCTIONS FOR TYPESETTING CAMERA-READY MANUSCRIPTS USING COMPUTER SOFTWARE^{∗}

Myname Andmylastname

October 10, 2002

Abstract

The abstract should summarize the context, content and conclusions of the paper in less than 200 words. It should not contain any references or displayed equations. Typeset the abstract in 8 pt Times Roman with baselineskip of 10 pt, making an indentation of 1.5 pica on the left and right margins. Typeset similarly for keywords below.^{1}

1 General Appearance

Contributions to the International Journal of Foundations of Computer Science will be reproduced by photographing the author’s submitted typeset manuscript. It is therefore essential that the manuscript be in its final form, and of good appearance because it will be printed directly without any editing. The manuscript should also be clean and unfolded. The copy should be evenly printed on a high resolution printer (300 dots/inch or higher). If typographical errors cannot be avoided, use cut and paste methods to correct them. Smudged copy, pencil or ink text corrections will not be accepted. Do not use cellophane or transparent tape on the surface as this interferes with the picture taken by the publisher’s camera.

∗For the title, try not to use more than 3 lines. Typeset the title in 10 pt Times Roman, uppercase and boldface.

1This is an abstract footnote

2 The Main Text

Contributions are to be in English. Authors are encouraged to have their contribution checked for grammar. American spelling should be used. Abbreviations are allowed but should be spelt out in full when first used. Integers ten and below are to be spelt out. Italicize foreign language phrases (e.g. Latin, French).

The text is to be typeset in 10 pt Times Roman, single spaced with baselineskip of 13 pt. Text area (excluding running title) is 5 inches (30 picas) across and 7.8 inches (47 picas) deep. Final pagination and insertion of running titles will be done by the publisher. Number each page of the manuscript lightly at the bottom with a blue pencil. Reading copies of the paper can be numbered using any legible means (typewritten or handwritten).

3 Major Headings

Major headings should be typeset in boldface with the first letter of important words capitalized.

3.1 Sub-headings

Sub-headings should be typeset in italic with the first letter of important words capitalized.

3.1.1 Sub-subheadings

Typeset sub-subheadings in medium face and capitalize the first letter of the first word only.

3.2 Numbering and Spacing

Sections, sub-sections and sub-subsections are numbered in Arabic. Use double spacing after major and subheadings, and single spacing after sub-subheadings.

3.3 Lists of Items

Lists may be laid out with each item marked by a dot:

one can define the following list formats to be newenvironments.

• item one,

• item two.

Items may also be numbered in lowercase Roman numerals:

1. item one

2. item two

4 Equations

Displayed equations should be numbered consecutively in each section, with the number set flush right and enclosed in parentheses.

μ(n, t) = \frac{∑_{i=1}^{∞} 1(d_{i} < t, N(d_{i}) = n)}{∫_{σ=0}^{t} 1(N(σ) = n)dσ}  (1)

Equations should be referred to in abbreviated form, e.g. “Eq.(1)” or “(1)”. In multiple-line equations, the number should be given on the last line.

Displayed equations are to be centered on the page width. Standard English letters like x are to appear as x (italicized) in the text if they are used as mathematical symbols. Punctuation marks are used at the end of equations as if they appeared directly in the text.

Theorem 1 Theorems, lemmas, etc. are to be numbered consecutively in the paper.

5 Illustrations and Photographs

Figures are to be inserted in the text nearest their first reference. Original india ink drawings of glossy prints are preferred. Please send one set of originals with copies. If the author requires the publisher to reduce the figures, ensure that the figures (including letterings and numbers) are large enough to be clearly seen after reduction. If photographs are to be used, only black and white ones are acceptable.

Figure 1: Labeled tree T.

Figures are to be sequentially numbered in Arabic numerals. The caption must be placed below the figure. Typeset in 8 pt Times Roman with baselineskip of 10 pt. Use double spacing between a caption and the text that follows immediately.

Previously published material must be accompanied by written permission from the author and publisher.

6 Tables

Tables should be inserted in the text as close to the point of reference as possible. Some space should be left above and below the table.

Tables should be numbered sequentially in the text in Arabic numerals. Captions are to be centralized above the tables. Typeset tables and captions in 8 pt Times Roman with baselineskip of 10 pt.

Table 1: Number of tests for WFF triple NA = 5, or NA = 8.

NP

3  5  8  10
3 1200 2000 2500 3000
NC 5 2000 2200 2700 3400
8 2500 2700 16000 22000
10 3000 3400 22000 28000

If tables need to extend over to a second page, the continuation of the table should be preceded by a caption, e.g. “Table 2. Cont’d.”

7 References

References in the text are to be numbered consecutively in Arabic numerals. They should be listed in alphabetical order of the first author (Use the style shown in the examples listed in References section). They are to be cited in bracketed form, as in this example [1, 2, 3]. Where it might cause ambiguity, cite references in brackets in abbreviated form, e.g. see Refs. [[1, 2]]. or Refs. [3, 6].

8 Footnotes

Footnotes should be numbered sequentially in superscript lowercase Roman letters^{2}.

2Footnotes should be typeset in 8 pt Times Roman at the bottom of the page.

9 Acknowledgements

This section should come before the References. Funding information may also be included here.

References

[1] M. J. Beeson, Foundations of Constructive Mathematics (Springer, Berlin, 1985).

[2] K. L. Clark, “Negations as failure,” in Logic and Data Bases, eds. H. Gallaire and J. Winker (Plenum Press, New York, 1973) pp. 293–306.

[3] D. Dolve, “Unanimity in an unknown and unreliable environment,” Proc. 22nd Annual Symposium on Foundations of Computer Science, Nashville, TN, Oct. 1981, pp. 159–168.

[4] W. L. Gewirtz, “Investigations in the theory of descriptive complexity,” Ph. D. Thesis, New York University, 1974.

[5] M. Joliat, “A simple technique for partial elimination of unit productions from LR(k) parsers,” IEEE Trans. Comput. C-27 (1976) 753–764.

[6] R. Lorentz and D. B. Benson, “Deterministic and nondeterministic flowchart interpretations,” J. Comput. System Sci. 27 (1983) 400–433.

[7] R. Tamassia, C. Batini and M. Talamo, “An algorithm for automatic layout of entity relationship diagrams,” in Entity-Relationship Approach to Software Engineering, Proc. 3rd Int. Conf. on Entity-Relationship Approach, eds. C. G. Davis, S. Jajodia, P. A. Ng and R. T. Yeh (North-Holland, Amsterdam, 1983) pp. 421–439.

Appendices should be used only when absolutely necessary. They should come after the References. If there is more than one appendix, number them alphabetically. Number displayed equations occurring in the Appendix in this way, e.g. (A.1), (A.2), etc.